OpenStudy (anonymous):

Bob and Joy Salkind want to save $50,000 in 5 ½ years for home improvement projects. If the Bank of Aventura is paying 8% interest compounded quarterly, how much must the couple deposit now to have the money for the projects?

OpenStudy (anonymous):

P(1+.08/4)^(4*5.5) = 50000 ??

OpenStudy (wolf1728):

First we should get the annual rate for 8% compounded quarterly annual % = [(1 + Rate / compounding periods)^comp periods]-1 annual % = [(1.02)^4]-1 annual % = .08243216 or 8.243216 To find the total, the formula is Total = Principal × (1 + Rate)^years So we can say Principal = Total / (1+Rate)^years Principal = 50,000 / (1.08243216)^5.5 Principal = 32,341.95